OpenStudy (kaylak12345):

Let g(x) be the reflection of f(x) = x² + 5 in the x-axis. What is a function rule for g(x)?

OpenStudy (mathmale):

You might want to begin with an easier example. Start with y=x^2. Reflect its graph in the x-axis. What would be the formula for this reflected graph?

OpenStudy (kaylak12345):

-x^2?

OpenStudy (kaylak12345):

@triciaal

OpenStudy (mathmale):

Yes, y=-x^2 is the function representing the reflection of y=x^2 in the x-axis. Now, going back to your original question, can you find the appropriate function rule?

OpenStudy (kaylak12345):

the answer is y=-x^2-5?

OpenStudy (triciaal):